### v1

# A system of NLS arising in optical material without Galilean symmetry

Abstract
We consider a system of NLS with cubic interactions arising in nonlinear optics without Galilean symmetry. The absence of Galilean symmetry can lead to many difficulties, such as global existence and blowup problems; see [Comm. Partial Differential Equations 46, 11 (2021), 2134–2170]. In this paper, we mainly focus on the influence of the absence of this symmetry on the traveling waves of the NLS system. Firstly, we obtain the existence of traveling solitary wave solutions that are non-radial and complex-valued. Secondly, using the asymptotic analysis method, when the frequency $\omega$ is sufficiently large, we establish the high frequency limit of the traveling solitary wave solution. Finally, for the mass critical case, we provide a novel condition for the existence of global solutions which is significantly different from the classical. In particular, this new condition breaks the traditional optimal assumption about initial data.
Keywords: Schrödinger system; Galilean symmetry; Traveling solitary waves; High frequency limit; Global existence
2020 Mathematics Subject Classification: Primary 35Q55; Secondary 35C07, 35B40
1 Introduction and Main Results
In this paper, we consider the Cauchy problem for the following system of nonlinear Schrödinger equations with cubic interaction
$\begin{cases}i\partial_{t}u+\Delta u+\left(\frac{1}{9}|u|^{2}+2|v|^{2}\right)u+\frac{1}{3}\bar{u}^{2}v=0,\\ i\gamma\partial_{t}v+\Delta v+(9|v|^{2}+2|u|^{2})v+\frac{1}{9}u^{3}=0,\end{cases}$ (1.1)
with the initial datum $(u,v)|_{t=0}=(u_{0},v_{0})$. Here $u,v:\mathbb{R}\times\mathbb{R}^{N}\to\mathbb{C}$ and the parameter $\gamma>0$, $\bar{u}$ is the complex conjugate of $u$.
System (1.1) arises from a system of nonlinear Schrödinger equations in a suitable dimensionless form, see [21], where the resonant interaction between the beam of the fundamental frequency and its third harmonic in a diffractive dielectric waveguide is characterized by one transverse dimension. For more details about the physical meaning of this system, one can refer to [5, 22] and the references therein. Models such as (1.1) so-called cascading nonlinear processes. These processes can indeed generate effective higher-order nonlinearities, and they stimulated the study of spatial solitary waves in optical materials with $\chi^{(2)}$ or $\chi^{(3)}$ susceptibilities. Recently, the study of system (1.1) attracted a quantity of attention; the topics cover existence, scattering, blowup, stability, etc. see [19, 18, 8, 1, 6].
Let us review some basic properties of system (1.1). It is well known that the Cauchy problem (1.1) is an infinite-dimensional Hamiltonian system and local well-posedness in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$ (see [3]), which have the following three conservation laws:
Mass:
$M_{3\gamma}(u(t),v(t))=\|u\|_{L^{2}}^{2}+3\gamma\|v\|_{L^{2}}^{2}=M_{3\gamma}(u_{0},v_{0});$ (1.2)
Energy:
$E(u(t),v(t))=\frac{1}{2}K(u(t),v(t))-D(u(t),v(t))=E(u_{0},v_{0}),$ (1.3)
where
$K(f,g)=\|\nabla f\|_{L^{2}}^{2}+\|\nabla g\|_{L^{2}}^{2},$
$D(f,g)=\int\frac{1}{36}|f(x)|^{4}+\frac{9}{4}|g(x)|^{4}+|f(x)|^{2}|g(x)|^{2}+\frac{1}{9}\Re\left(\bar{f}^{3}(x)g(x)\right);$ (1.4)
Momentum:
$P(u(t),v(t))=(i\nabla u(t),u(t))+\gamma(i\nabla v(t),v(t))=P(u_{0},v_{0}),$ (1.5)
where
$(u,v)=\Re\int u\bar{v}.$
Notice that for $\gamma=3$, which is called the mass resonance condition, system (1.1) is invariant under the Galilean transformation
$(u,v)\mapsto\left(e^{\frac{i}{2}c\cdot x-\frac{i}{4}|c|^{2}t}u(t,x-ct),e^{\frac{i\gamma}{2}c\cdot x-\frac{i\gamma}{4}|c|^{2}t}v(t,x-ct)\right).$
for any $c\in\mathbb{R}^{N}$. If $\gamma\neq 3$, the system (1.1) is not invariant under the Galilean transformations.
In this paper, we are interested in the traveling solitary wave solution of system (1.1). Now, we consider traveling solitary wave solutions of (1.1) in the form of
$\left(u_{c,w}(t,x),v_{c,w}(t,x)\right)=\left(e^{i\omega t}\phi(x-ct),e^{3i\omega t}\psi(x-ct)\right).$ (1.6)
In particular, when $c=0$, (1.6) is the form of standing wave solutions, then the system (1.1) reads
$\begin{cases}-\Delta\phi+\omega u=\left(\frac{1}{9}\phi^{2}+2\psi^{2}\right)\phi+\frac{1}{3}\bar{\phi}^{2}\psi,\\ -\Delta\psi+3\gamma\omega u=\left(9\psi^{2}+2\phi^{2}\right)\psi+\frac{1}{9}\phi^{3}.\end{cases}$ (1.7)
The existence of the ground state was obtained by Oliveira and Pastor, see [18]. When $\gamma=3$, one can obtain traveling solitary wave solutions from the standing wave solutions through the Galilean transformation. However, when $\gamma\neq 3$, such a construction does not work due to the lack of Galilean symmetry, Therefore, it is meaningful to study the existence of traveling solitary wave solutions of (1.1) when $\gamma\neq 3$.
The question of the existence and properties of the traveling solitary wave solution of the system (1.1) is widely open. However, in other settings, like the general nonlinear Schrödinger system
$\begin{cases}i\partial_{t}u+\Delta u+\left(a_{1}|u|^{2}+b|v|^{2}\right)u=0,\\ i\partial_{t}v+\Delta v+(a_{2}|v|^{2}+b|u|^{2})v=0,\end{cases}$ (1.8)
standing wave solutions are studied by the variational method. There is a very extensive literature on this subject; we refer, for example, to [24, 23, 20, 14, 16, 4] and references therein. Due to the system (1.8) is invariant under the Galilean transformation. This suggests that one can obtain the desired result by considering the real elliptic system.
However, in this paper, we will focus on the complex elliptic system. For $(\omega,c)\in\mathbb{R}\times\mathbb{R}^{N}$ and $c\neq 0$, (1.6) is a solution of (1.1) if and only if
$\left(\phi(x-ct),\psi(x-ct)\right)$
is a solution of the stationary system
$\begin{cases}-\Delta\phi+\omega\phi+ic\cdot\nabla\phi-\left(\frac{1}{9}|\phi|^{2}+2|\psi|^{2}\right)\phi-\frac{1}{3}\bar{\phi}^{2}\psi=0,\\ -\Delta\psi+3\gamma\omega\psi+i\gamma c\cdot\nabla\psi-\left(9|\psi|^{2}+2|\phi|^{2}\right)\psi-\frac{1}{9}\phi^{3}=0.\end{cases}$ (1.9)
The energy functional with respect to (1.9) is defined by
$S_{\omega,c}(u,v)=$ $E(u,v)+\frac{1}{2}\omega M_{3\gamma}(u,v)+\frac{1}{2}c\cdot P(u,v),$ (1.10)
where $E$, $M_{3\gamma}$ and $P$ are given by (1.3), (1.2) and (1.5), respectively. For the sake of argument, we rewrite the energy functional (1.10) as
$S_{\omega,c}(u,v)=$ $\frac{1}{2}\int\left|\nabla\left(e^{-\frac{i}{2}c\cdot x}u\right)\right|^{2}+\frac{1}{2}\left(\omega-\frac{|c|^{2}}{4}\right)\|u\|_{L^{2}}^{2}$
$+\frac{1}{2}\int\left|\nabla\left(e^{-\frac{i\gamma}{2}c\cdot x}v\right)\right|^{2}+\frac{1}{2}\left(3\gamma\omega-\frac{\gamma^{2}|c|^{2}}{4}\right)\|v\|_{L^{2}}^{2}-D(u,v).$
Now we define the function space $X_{\omega,c}$ by
$X_{\omega,c}=\left\{(u,v):\left(e^{-\frac{i}{2}c\cdot x}u,e^{-\frac{i\gamma}{2}c\cdot x}v\right)\in\tilde{X}_{\omega,c}\right\},$
where
$\tilde{X}_{\omega,c}=H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N}),$
with the norm
$\|(u,v)\|_{\tilde{X}_{\omega,c}}=\|\nabla u\|_{L^{2}}^{2}+\|\nabla v\|_{L^{2}}^{2}+\omega\|u\|_{L^{2}}^{2}+3\gamma\omega\|v\|_{L^{2}},~{}\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\},~{}\gamma>0.$
Notice that $S_{\omega,c}$ is defined on $X_{\omega,c}$ if
$0\displaystyle 1\leq N\leq 3,~{}\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\},~{}\gamma>$ 0 . (1.11)
This condition comes from the Sobolev embedding, which is used to control the nonlinear term. In this paper, we consider the case $\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$ and $N\leq 3$.
We denote the set of all non-zero solutions of (1.9) by
$\mathcal{A}_{\omega,c}=\{(\phi,\psi)\in X_{\omega,c}~{}:~{}(\phi,\psi)\neq(0,0),~{}S^{\prime}_{\omega,c}(\phi,\psi)=0\}$
and the set of all boosted ground states
$\mathcal{G}_{\omega,c}=\{(\phi,\psi)\in\mathcal{A}_{\omega,c},~{}S_{\omega,c}(\phi,\psi)\leq S_{\omega,c}(\phi_{1},\psi_{1}),~{}\text{for all}~{}(\phi_{1},\psi_{1})\in\mathcal{A}_{\omega,c}\}.$
In particular, if $c=0$, $\mathcal{G}_{\omega,0}$ is the set of all ground states.
For $(\phi,\psi)$ satisfying (1.9), let $(\phi,\psi)=\left(e^{\frac{i}{2}c\cdot x}\tilde{\phi},e^{\frac{i\gamma}{2}c\cdot x}\tilde{\psi}\right)$. Then $(\tilde{\phi},\tilde{\psi})$ satisfies
$\begin{cases}-\Delta\tilde{\phi}+\left(\omega-\frac{|c|^{2}}{4}\right)\tilde{\phi}-\left(\frac{1}{9}|\tilde{\phi}|^{2}+2|\tilde{\psi}|^{2}\right)\tilde{\phi}-\frac{1}{3}e^{i\left(\frac{\gamma}{2}-\frac{3}{2}\right)c\cdot x}\bar{\tilde{\phi}}^{2}\tilde{\psi}=0,\\ -\Delta\tilde{\psi}+\left(3\gamma\omega-\frac{|c|^{2}}{4}\right)\tilde{\psi}-\left(9|\tilde{\psi}|^{2}+2|\tilde{\phi}|^{2}\right)\tilde{\psi}-\frac{1}{9}e^{i\left(\frac{3}{2}-\frac{\gamma}{2}\right)c\cdot x}\tilde{\phi}^{3}=0.\end{cases}$ (1.12)
Now we state our first result.
Theorem 1.1.
Let $1\leq N\leq 3$, $\gamma>0$, $\gamma\neq 3$ and $\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$.
(i) Then there exists a boosted ground state solution $(u,v)$ of (1.9) with $v\neq 0$. That is, the system (1.1) has a traveling solitary wave solution in the form of (1.6).
(ii) If $3\gamma>\left(\frac{9}{\sqrt{2}}\right)^{\frac{4}{N-4}}$, then the traveling solitary wave solutions obtained in (i) are non-trivial, that is, $u\neq 0$ and $v\neq 0$.
Theorem 1.2.
Let $1\leq N\leq 3$ and $\gamma>0$. Then there is not a non-zero traveling solitary wave solution in the form of (1.6) provided one of the following conditions:
(i) $N=2$ and $\omega\leq 0$.
(ii) $N=1$ or $N=3$, $\omega\leq-\frac{|c|}{2N}$ with $|c|<\min\left\{2(4-N),\frac{6}{\gamma}(4-N)\right\}$.
Comments:
1. The existence of traveling solitary wave solutions for (1.1) is equivalent to the existence of solutions for (1.12). Note that the non-trivial solutions of (1.12) with $\gamma\neq 3$ (non-mass resonance condition) are essentially non-radial and complex-valued. Therefore, the lack of symmetries yields the new and non-trivial existence result.
2. In Theorem 1.1 (i), the traveling solitary wave solution may be semi-trivial or non-trivial. That is, there exists a semi-trivial solution $(0,\tilde{Q})$, where $\tilde{Q}$ is a Galilean transformation of $Q$ and $Q$ is the positive ground state of
$-\Delta Q+3\gamma\omega Q=9Q^{3}.$ (1.13)
This is completely different from the case $\chi^{(2)}$. In optical material with $\chi^{(2)}$, the system read as
$\begin{cases}i\partial_{t}u+\frac{1}{2m}\Delta u=\bar{u}v,\\ i\partial_{t}v+\frac{1}{2M}\Delta u=u^{2},\end{cases}(t,x)\in\mathbb{R}\times\mathbb{R}^{N},~{}~{}m,M>0.$ (1.14)
Fukaya, Hayashi and Inui [9] obtained the existence of traveling solitary wave solutions without mass resonance. In addition, the first author in [12] obtained the similar result for three-wave interaction. In these two cases, the system does not have semi-trivial traveling solitary wave solutions, so the argument is simple. For more about the system with quadratic interaction like (1.14), one can see [7, 11, 17, 10] and the reference therein.
Remark 1.3.
1. When $0<3\gamma\leq\left(\frac{9}{\sqrt{2}}\right)^{\frac{4}{N-4}}$ and $\gamma\neq 3$, the traveling solitary wave solutions may be semi-trivial or non-trivial. This case is left as an interesting open problem.
2. In Theorem 1.1, we assume that $\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$ and obtain the existence of traveling solitary waves. In Theorem 1.2, we assume that $\omega\leq 0$ or $\omega\leq-\frac{|c|}{2N}$ and obtain the non-existence of traveling solitary waves. However, for $0<\omega<\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$, $N=2$ or $\frac{|c|}{2N}<\omega<\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$, $N=1,3$, the existence or non-existence result is left as an interesting open problem.
Next, we aim to study the high frequency limit of the boosted ground state of (1.1).
Theorem 1.4.
Let $1\leq N\leq 3$ and $(u_{\omega},v_{\omega})$ be the nontrivial traveling solitary wave solution of (1.9) (see Theorem 1.1). Assume that $|c|$ is bounded and $\omega\to+\infty$, then we have
[…]
Here $(u_{\infty},v_{\infty})$ is the complex ground state solution of the system (1.9) with $c=0$ and $\omega=1$.
Comments:
1. In Theorem 1.4, we consider the high frequency limit when the velocity $|c|$ is bounded and the frequency $\omega\to+\infty$. In fact, we can also consider the situation when $|c|\to 0$ and $\omega$ is bounded.
2. Unlike the single equation, see [2, 15], there is no momentum term $P(u,v)$, and it is relatively easy to handle. In our case, things become more difficult since we do not know the sign of the term $c\cdot P(u,v)$ and the functions are all complex-valued.
Finally, we provide an interesting result. This result is the global existence of (1.1) for the $L^{2}$-critical case $(N=2)$. When $N=2$, from [18], it is known that if the initial data $(u_{0},v_{0})\in H^{1}(\mathbb{R}^{2})\times H^{1}(\mathbb{R}^{2})$ and
$M_{3\gamma}(u_{0},v_{0})<M_{3\gamma}(Q,P),$ (1.15)
then the solution of (1.1) is globally, where $(Q,P)$ is the positive ground state solution of the following system:
$\begin{cases}-\Delta u+u=\frac{1}{9}u^{3}+2v^{2}u+\frac{1}{3}u^{2}v,\\ -\Delta v+3\gamma v=9v^{3}+2u^{2}v+\frac{1}{9}u^{3}.\end{cases}$
From the sharp Gagliardo-Nirenberg inequality
$D(u,v)\leq C_{opt}^{(1)}K(u,v)M_{3\gamma}(u,v),~{}~{}(u,v)\in H^{1}(\mathbb{R}^{2})\times H^{1}(\mathbb{R}^{2}),$ (1.16)
where
$C_{opt}^{(1)}=\frac{1}{2}\frac{1}{M_{3\gamma}(Q,P)}.$
We know that the condition (1.15) is sharp in general. Indeed, there exist $u_{0},v_{0}\in H^{1}(\mathbb{R}^{2})$ satisfying
$M_{3\gamma}(u_{0},v_{0})=M_{3\gamma}(Q,P),$
such that the system (1.1) has the blowup solution, see [18, Theorem 4.2]. It was also proved that there exists a blowup solution in the weight space $L^{2}(\mathbb{R}^{2},|x|^{2}dx)$, see [18, Theorem 4.7 and Theorem 4.8].
On the other hand, from Appendix B, we have the following sharp Gagliardo-Nirenberg inequality, for $(u,v)\in H^{1}(\mathbb{R}^{2})\times H^{1}(\mathbb{R}^{2})$,
$\int\frac{1}{36}|u|^{4}+\frac{9}{4}|v|^{4}+|u|^{2}|v|^{2}\leq C_{opt}^{(2)}K(u,v)M_{3\gamma}(u,v),$ (1.17)
where
$C_{opt}^{(2)}=\frac{1}{2}\frac{1}{M_{3\gamma}(Q^{*},P^{*})}.$
Here $(Q^{*},P^{*})$ is the positive ground state of the following system
$\begin{cases}-\Delta u+u=\frac{1}{9}u^{3}+2v^{2}u,\\ -\Delta v+3\gamma v=9v^{3}+2u^{2}v.\end{cases}$ (1.18)
Obviously, from (1.16) and (1.17), we have $C_{opt}^{(1)}>C_{opt}^{(2)}$, this means that
$M_{3\gamma}(Q,P)<M_{3\gamma}(Q^{*},P^{*}).$
Now we give the following global result.
Theorem 1.5.
Let $N=2$, $c\in\mathbb{R}^{2}$, $u_{0},v_{0}\in H^{1}(\mathbb{R}^{2})\backslash\{0\}$ and
$M_{3\gamma}(u_{0},v_{0})<M_{3\gamma}(Q^{*},P^{*}).$ (1.19)
Set
$(u_{0,c},v_{0,c})=\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right).$ (1.20)
Then the following statements hold.
(i) If $\gamma<3$, then there exist $A_{0},A_{1}>0$ such that $\|v_{0}\|_{L^{2}}^{2}<A_{0}$ and $|c|>A_{1}$, the $H^{1}$ solution of (1.1) with the initial data $(u(0),v(0))=(u_{0,c},v_{0,c})$ exists globally in time.
(ii) If $\gamma>3$, then there exist $B_{0},B_{1}>0$ such that $\|u_{0}\|_{L^{2}}^{2}<B_{0}$ and $|c|>B_{1}$, the $H^{1}$ solution of (1.1) with the initial data $(u(0),v(0))=(u_{0,c},v_{0,c})$ exists globally in time.
Comments:
1. Using mass, energy conservation and the Gagliardo-Nirenberg inequality, one can easily obtain the global existence with the initial data $(u_{0},v_{0})\in H^{1}(\mathbb{R})\times H^{1}(\mathbb{R})$ in one dimensional case. For three dimensional case, we only obtain the global existence in the set $A_{\omega,c}^{+}$, see (4.1) and Lemma 4.2, this is similar to the classical case.
2. From [18], the condition (1.15) is sharp. But in our result, we give a new condition: (1.19) and the restriction one of the initial data $\|u_{0}\|_{L^{2}}$ or $\|v_{0}\|_{L^{2}}$, for the existence of global solution. This condition is weaker than (1.15) since $M_{3\gamma}(Q,P)<M_{3\gamma}(Q^{*},P^{*})$. This new phenomenon means that when $\gamma\neq 3$, the momentum change of the initial data by (1.20) essentially influences the global properties of the solution, which comes from the lack of Galilean invariance.
3. This is different from the global existence in [9, Theorem 1.3], where they only restrict one of the $u(0)$ or $v(0)$. In fact, in [9], they considered the nonlinear NLS system with quadratic interaction, but in our case, we must add the condition (1.19), which comes from the cubic interaction.
This paper is organized as follows:. In Section 2, we study the variational problem and show the existence and non-existence of traveling solitary wave solutions (Theorem 1.1 and 1.2). In Section 3, we establish the high frequency limit of the traveling solitary waves. In Section 4, we establish the global existence (Theorem 1.5). The final section is an appendix.
2 Existence and non-existence of traveling solitary waves
In this section, we aim to prove the existence and non-existence of traveling solitary wave solutions by solving variational problems on the Nehari manifold. It is worth noting that we will be dealing with a complex elliptic system.
For the sake of simplicity, we use the following notation. We set
$Q_{\omega,c}(u,v)=$ $\frac{1}{2}\int|\nabla u|^{2}+\frac{1}{2}\int|\nabla v|^{2}+\frac{\omega}{2}\|u\|_{L^{2}}^{2}+\frac{3\omega\gamma}{2}\|v\|_{L^{2}}^{2}+\frac{1}{2}c\cdot P(u,v).$ (2.1)
The energy functional (1.10) reads
$S_{\omega,c}(u,v)=Q_{\omega,c}(u,v)-D(u,v),$
where $D$ is defined in (1.4). Now we introduce the Nehari functional
$N_{\omega,c}(u,v)=\partial_{\tau}S_{\omega,c}(\tau u,\tau v)|_{\tau=1}=2Q_{\omega,c}(u,v)-4D(u,v).$ (2.2)
The minimization problem
$\mu_{\omega,c}=\inf\{S_{\omega,c}(\phi,\psi)~{}:~{}(\phi,\psi)\in\mathcal{N}_{\omega,c}\},$ (2.3)
where
$\mathcal{N}_{\omega,c}=\{(\phi,\psi)\in X_{\omega,c},~{}(\phi,\psi)\neq(0,0),~{}N_{\omega,c}(\phi,\psi)=0\}.$
We define the minimizers $a_{\omega,c}$ by
$a_{\omega,c}=\{(\phi,\psi)\in\mathcal{N}_{\omega,c}~{}:~{}S_{\omega,c}(\phi,\psi)=\mu_{\omega,c}\}.$
We also let
$\widetilde{Q}_{\omega,c}(u,v)=$ $\frac{1}{2}\int|\nabla u|^{2}+\frac{1}{2}\int|\nabla v|^{2}+\frac{1}{2}\left(\omega-\frac{|c|^{2}}{4}\right)\|u\|_{L^{2}}^{2}+\frac{1}{2}\left(3\gamma\omega-\frac{\gamma^{2}|c|^{2}}{4}\right)\|v\|_{L^{2}}^{2},$
$\widetilde{D}(u,v)=$ $\int\left(\frac{1}{36}|u(x)|^{4}+\frac{9}{4}|v(x)|^{4}+|u(x)|^{2}|v(x)|^{2}+\frac{1}{9}\Re\left[e^{\frac{i(3-\gamma)}{2}c\cdot x}\left(\bar{u}^{3}(x)v(x)\right)\right]\right).$
If
$(\tilde{u},\tilde{v})=\left(e^{-\frac{i}{2}c\cdot x}u,e^{-\frac{i\gamma}{2}c\cdot x}v\right),$
then we have the following relations
$\tilde{Q}_{\omega,c}(\tilde{u},\tilde{v})=Q_{\omega,c}(u,v),~{}~{}\widetilde{D}(\tilde{u},\tilde{v})=D(u,v).$
The corresponding functionals
$\tilde{S}_{\omega,c}(u,v)=$ $\tilde{Q}_{\omega,c}(u,v)-\tilde{D}(u,v),$
$\tilde{N}_{\omega,c}(u,v)=$ $\partial_{\tau}\tilde{S}_{\omega,c}(\tau u,\tau v)|_{\tau=1}=2\tilde{Q}_{\omega,c}(u,v)-4\tilde{D}(u,v).$
We also use the following notations.
$\widetilde{\mathcal{A}}_{\omega,c}=\{(\phi,\psi)\in\tilde{X}_{\omega,c}~{}:~{}(\phi,\psi)\neq(0,0),~{}\tilde{S}^{\prime}_{\omega,c}(\phi,\psi)=0\},$
$\widetilde{\mathcal{G}}_{\omega,c}=\{(\phi,\psi)\in\widetilde{\mathcal{A}}_{\omega,c},~{}\tilde{S}_{\omega,c}(\phi,\psi)\leq\tilde{S}_{\omega,c}(\phi_{1},\psi_{1}),~{}\text{for all}~{}(\phi_{1},\psi_{1})\in\tilde{\mathcal{A}}_{\omega,c}\},$
$\tilde{\mu}_{\omega,c}=\inf\{\tilde{S}_{\omega,c}(\phi,\psi)~{}:~{}(\phi,\psi)\in\tilde{\mathcal{N}}_{\omega,c}\},$
$\tilde{\mathcal{N}}_{\omega,c}=\{(\phi,\psi)\in\tilde{X}_{\omega,c},~{}(\phi,\psi)\neq(0,0),~{}\tilde{N}_{\omega,c}(\phi,\psi)=0\},$
$\tilde{a}_{\omega,c}=\{(\phi,\psi)\in\tilde{\mathcal{N}}_{\omega,c}~{}:~{}\tilde{S}_{\omega,c}(\phi,\psi)=\tilde{\mu}_{\omega,c}\}.$
Notice that
$(\phi,\psi)\in{\mathcal{G}}_{\omega,c}\Leftrightarrow(\tilde{\phi},\tilde{\psi})\in\tilde{\mathcal{G}}_{\omega,c},~{}~{}(\phi,\psi)\in a_{\omega,c}\Leftrightarrow(\tilde{\phi},\tilde{\psi})\in\tilde{a}_{\omega,c},$
and
$\mu_{\omega,c}=\tilde{\mu}_{\omega,c}.$
Lemma 2.1.
Let $(\phi,\psi)$ be the solution of (1.9), then we have the following identities:
$\int\left(|\nabla\phi|^{2}+|\nabla\psi|^{2}\right)+\omega\int\left(|\phi|^{2}+3\gamma|\psi|^{2}\right)+(ic\cdot\nabla\phi,\phi)+(i\gamma c\cdot\nabla\psi,\psi)=4D(\phi,\psi).$ (2.4)
And
$\frac{N-2}{2}\int\left(|\nabla\phi|^{2}+|\nabla\psi|^{2}\right)+\frac{N}{2}\int\left(\omega|\phi|^{2}+3\gamma\omega|\psi|^{2}\right)$
$+\frac{N-1}{2}\left[(\phi,ic\cdot\nabla\phi)+\gamma(\psi,ic\cdot\nabla\psi)\right]=ND(\phi,\psi).$ (2.5)
For the readers’ convenience, we give the proof in Appendix A.
Now we state the main result in this section.
Proposition 2.2.
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$, then
$\mathcal{G}_{\omega,c}=a_{\omega,c}\neq\emptyset.$
In fact, this is equivalent to $\widetilde{\mathcal{G}}_{\omega,c}=\tilde{a}_{\omega,c}\neq\emptyset.$
Remark 2.3.
If $(u,v)\in\mathcal{G}_{\omega,c}$, then $v\neq 0$. In fact, if $v=0$, from system (1.9), one can easily obtain that $u=0$.
If Proposition 2.2 is true, then we can easily obtain the Theorem 1.1 (i). In order to prove Proposition 2.2, we need the following lemmas.
Lemma 2.4.
Let $\gamma\neq 3$, then $\tilde{a}_{\omega,c}\subset\widetilde{\mathcal{G}}_{\omega,c}$.
Proof.
Let $(\phi,\psi)\in\tilde{a}_{\omega,c}$. Since $\tilde{N}_{\omega,c}(\phi,\psi)=0$ and $(\phi,\psi)\neq 0$, we have
$\left(\tilde{N}^{\prime}_{\omega,c}(\phi,\psi),(\phi,\psi)\right)=4\tilde{Q}_{\omega,c}(\phi,\psi)-16\tilde{D}(\phi,\psi)=-4\tilde{Q}_{\omega,c}(\phi,\psi)<0.$ (2.6)
By the Lagrange multiplier theorem, there exists $\lambda\in\mathbb{R}$ such that $\tilde{S}^{\prime}_{\omega,c}(\phi,\psi)=\lambda\tilde{N}^{\prime}_{\omega,c}(\phi,\psi)$. Moreover, we have
$\lambda\left(\tilde{N}^{\prime}_{\omega,c}(\phi,\psi),(\phi,\psi)\right)=\left(\tilde{S}^{\prime}_{\omega,c}(\phi,\psi),(\phi,\psi)\right)=\tilde{N}_{\omega,c}(\phi,\psi)=0.$ (2.7)
Combining (2.6) and (2.7), we can obtain $\lambda=0$. Hence, $\tilde{S}^{\prime}_{\omega,c}(\phi,\psi)=0$, which implies that $(\phi,\psi)\in\tilde{\mathcal{A}}_{\omega,c}$.
Notice that $\tilde{\mathcal{A}}_{\omega,c}\subset\tilde{\mathcal{N}}_{\omega,c}$ and $(\phi,\psi)\in\tilde{a}_{\omega,c}$, we have
$\tilde{S}_{\omega,c}(\phi,\psi)\leq\tilde{S}_{\omega,c}(\phi_{1},\psi_{1})~{}~{}\text{for all}~{}~{}(\phi_{1},\psi_{1})\in\tilde{\mathcal{A}}_{\omega,c},$
which implies $(\phi,\psi)\in\tilde{\mathcal{G}}_{\omega,c}$. ∎
Lemma 2.5.
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. Assume that $\tilde{a}_{\omega,c}\neq\emptyset$, then $\widetilde{\mathcal{G}}_{\omega,c}\subset\tilde{a}_{\omega,c}$.
Proof.
Let $(\phi,\psi)\in\widetilde{\mathcal{G}}_{\omega,c}$. One can take $(\phi_{1},\psi_{1})\in\tilde{a}_{\omega,c}$, by Lemma 2.4, we have $(\phi_{1},\psi_{1})\in\widetilde{\mathcal{G}}_{\omega,c}$, i.e., $\tilde{S}_{\omega,c}(\phi_{1},\psi_{1})=\tilde{S}_{\omega,c}(\phi,\psi)$. Therefore, for each $(u,v)\in\tilde{\mathcal{N}}_{\omega,c}$, we obtain
$\tilde{S}_{\omega,c}(\phi,\psi)=\tilde{S}_{\omega,c}(\phi_{1},\psi_{1})\leq\tilde{S}_{\omega,c}(u,v).$
Since $(\phi,\psi)\in\widetilde{\mathcal{G}}_{\omega,c}\subset\tilde{\mathcal{N}}_{\omega,c}$, we deduce that $(\phi,\psi)\in\tilde{a}_{\omega,c}$. ∎
Lemma 2.6.
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. Then $\tilde{\mu}_{\omega,c}>0$.
Proof.
By the definition of $\tilde{S}_{\omega,c}$, $\tilde{Q}_{\omega,c}$ and $\tilde{N}_{\omega,c}$, we have the following relation
$\tilde{S}_{\omega,c}=\frac{1}{2}\tilde{Q}_{\omega,c}+\frac{1}{4}\tilde{N}_{\omega,c}.$
One can rewrite $\tilde{\mu}_{\omega,c}$ as
$\tilde{\mu}_{\omega,c}=\inf\left\{\frac{1}{2}\tilde{Q}_{\omega,c},~{}~{}(\phi,\psi)\in\tilde{\mathcal{N}}_{\omega,c}\right\}.$ (2.8)
By Hölder inequality, Gagliardo-Nirenberg inequality and Sobolev embedding theorem, we have
$\tilde{Q}_{\omega,c}(u,v)=2\widetilde{D}(u,v)$
$=$ $\int\left(\frac{1}{36}|u(x)|^{4}+\frac{9}{4}|v(x)|^{4}+|u(x)|^{2}|v(x)|^{2}+\frac{1}{9}\Re\left[e^{\frac{i(3-\gamma)}{2}c\cdot x}\left(\bar{u}^{3}(x)v(x)\right)\right]\right)$
$\leq$ $C\|\nabla u\|_{L^{2}}^{N}\|u\|_{L^{2}}^{4-N}+C\|\nabla v\|_{L^{2}}^{N}\|v\|_{L^{2}}^{4-N}+C\|\nabla u\|_{L^{2}}^{\frac{3N}{4}}\|u\|_{L^{2}}^{\frac{3(4-N)}{4}}\|\nabla v\|_{L^{2}}^{\frac{N}{4}}\|v\|_{L^{2}}^{\frac{(4-N)}{4}}$
$\leq$ $C\tilde{Q}^{2}_{\omega,c}(u,v).$
Dividing both sides of the above by $\tilde{Q}_{\omega,c}>0$, we obtain the positive lower bound for $\tilde{Q}_{\omega,c}$. ∎
Lemma 2.7.
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If $(u,v)\in\tilde{X}_{\omega,c}$ satisfies $\tilde{N}_{\omega,c}(u,v)<0$, then $\frac{1}{2}\tilde{Q}_{\omega,c}>\tilde{\mu}_{\omega,c}$.
Proof.
If $\tilde{N}_{\omega,c}(u,v)<0$, then $02\tilde{D}(u,v)>\tilde{Q}_{\omega,c}(u,v)>$ 0. From this, we have
$\lambda_{0}:=\frac{\tilde{Q}_{\omega,c}(u,v)}{2\tilde{D}(u,v)}\in(0,1)$
and $\tilde{N}_{\omega,c}(\lambda_{0}u,\lambda_{0}v)=0$. By (2.8), we deduce
$\tilde{\mu}_{\omega,c}\leq\frac{1}{2}\tilde{Q}_{\omega,c}(\lambda_{0}u,\lambda_{0}v)=\frac{\lambda_{0}^{2}}{2}\tilde{Q}_{\omega,c}(u,v)<\frac{1}{2}\tilde{Q}_{\omega,c}(u,v).$
Now we complete the proof of this Lemma. ∎
Lemma 2.8.
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If the sequence $\{(u_{n},v_{n})\}$ weakly converges to $(u,v)\in\tilde{X}_{\omega,c}$, then
[…]
Proof.
By direct calculation, one can obtain this result. ∎
Note that $\tilde{Q}_{\omega,c}$ and $\tilde{D}$ are invariant under
$\tilde{\tau}_{y}(u,v):=\left(e^{-\frac{i}{2}c\cdot x}u(\cdot-y),e^{-\frac{i\gamma}{2}c\cdot x}v(\cdot-y)\right),$
that is,
$\tilde{Q}_{\omega,c}(\tilde{\tau}_{y}(u,v))=\tilde{Q}_{\omega,c}(u,v),~{}~{}\tilde{D}(\tilde{\tau}_{y}(u,v))=\tilde{D}(u,v),$
for all $y\in\mathbb{R}^{N}$.
Lemma 2.9.
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If a sequence $\{(u_{n},v_{n})\}$ in $\tilde{X}_{\omega,c}$ satisfies
[…]
for some positive constants $A_{1},A_{2}>0$, then there exist $\{y_{n}\}$ and $(u,v)\in\tilde{X}_{\omega,c}\setminus\{(0,0)\}$ such that $\{\tilde{\tau}_{y}(u_{n},v_{n})\}$ has a subsequence that weakly converges to $(u,v)$ in $\tilde{X}_{\omega,c}$.
Proof.
Since $\tilde{Q}_{\omega,c}(u_{n},v_{n})\to A_{1}$, we deduce that the sequence $\{(u_{n},v_{n})\}$ is bounded in $\tilde{X}_{\omega,c}$. Since $\tilde{D}(u_{n},v_{n})\to A_{2}$, we obtain
$\limsup_{n\to\infty}\|u_{n}\|_{L^{4}}+\limsup_{n\to\infty}\|v_{n}\|_{L^{4}}>0.$
Hence, by [13, Lemma 6], we can obtain the desired result. ∎
Lemma 2.10.
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If a sequence $\{(u_{n},v_{n})\}$ in $\tilde{X}_{\omega,c}$ satisfies
[…]
then there exist $\{y_{n}\}$ and $(u,v)\in\tilde{X}_{\omega,c}\setminus\{(0,0)\}$ such that $\{\tilde{\tau}_{y_{n}}(u_{n},v_{n})\}$ has a subsequence that converges to $(u,v)$ in $\tilde{X}_{\omega,c}$. In particular, $(u,v)\in\tilde{a}_{\omega,c}$.
Proof.
By assumptions, we have
$\frac{1}{2}\tilde{Q}_{\omega,c}(u_{n},v_{n})=\tilde{S}_{\omega,c}(u_{n},v_{n})-\frac{1}{4}\tilde{N}(u_{n},v_{n})\to\tilde{\mu}_{\omega,c},$
$\tilde{D}(u_{n},v_{n})=\tilde{S}_{\omega,c}(u_{n},v_{n})-\frac{1}{2}\tilde{N}_{\omega,c}(u_{n},v_{n})\to\tilde{\mu}_{\omega,c}.$
By Lemma 2.6 and Lemma 2.9, there exist $\{y_{n}\}$, $(u,v)\in\tilde{X}_{\omega,c}\setminus\{(0,0)\}$, and a subsequence of $\tilde{\tau}_{y_{n}}(u_{n},v_{n})$(still denoted by $\tilde{\tau}_{y_{n}}(u_{n},v_{n})$) such that $\tilde{\tau}_{y_{n}}(u_{n},v_{n})\rightharpoonup(u,v)$ weakly in $\tilde{X}_{\omega,c}$.
By the weak convergence of $\tilde{\tau}_{y_{n}}(u_{n},v_{n})$ and Lemma 2.8, we have
$\tilde{Q}_{\omega,c}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))-\tilde{Q}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)\to\tilde{Q}_{\omega,c}(u,v),$ (2.9)
$\tilde{N}_{\omega,c}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))-\tilde{N}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)\to\tilde{N}_{\omega,c}(u,v).$ (2.10)
From (2.9) and $\tilde{Q}_{\omega,c}(u,v)>0$, we obtain that, up to subsequence,
$\frac{1}{2}\lim_{n\to\infty}\tilde{Q}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)<\frac{1}{2}\lim_{n\to\infty}\tilde{Q}_{\omega,c}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))=\tilde{\mu}_{\omega,c}.$
From this and Lemma 2.7, we obtain $\tilde{N}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)>0$ for $n$ large enough. Therefore, from (2.10), $\tilde{N}_{\omega,c}(u,v)\leq 0$ since $\tilde{N}_{\omega,c}(\tilde{\tau}_{y_{n}}(u_{n},v_{n}))\to 0$. Again, by Lemma 2.7 and the weakly semi-continuity of norms, we have
$\tilde{\mu}_{\omega,c}\leq\frac{1}{2}\tilde{Q}_{\omega,c}(u,v)\leq\frac{1}{2}\tilde{Q}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})\right)=\tilde{\mu}_{\omega,c}.$
Therefore, by (2.9), we deduce $\tilde{Q}_{\omega,c}\left(\tilde{\tau}_{y_{n}}(u_{n},v_{n})-(u,v)\right)\to 0$ as $n\to\infty$, which implies that $\tilde{\tau}_{y_{n}}(u_{n},v_{n})\to(u,v)$ strongly in $\tilde{X}_{\omega,c}$. Now we have completed the proof of this lemma. ∎
Proof of Proposition 2.2..
Combining the Lemmas 2.4, 2.5 and 2.10, one can get Proposition 2.2. ∎
In order to prove Theorem 1.1 (ii), we will prove that the minimizer $(u,v)$ is a non-semitrivial solution ($u\neq 0$ and $v\neq 0$) of the system (1.9). If $u=0$, one can take $v=\tilde{Q}$, where $\tilde{Q}$ is a Galilean transformation of $Q$, and $Q$ is the positive ground state of equation (1.13), then $(0,\tilde{Q})$ is the semi-trivial solution of system (1.9). Since $Q$ is the ground state of (1.13), then $S_{\omega,0}(Q)\leq S_{\omega,c}(\tilde{Q})$. As below, we give the relation between $S_{\omega,c}(u,v)$ and $S_{\omega,0}(0,Q)$.
Lemma 2.11.
Assume that $3\gamma>\left(\frac{9}{\sqrt{2}}\right)^{\frac{4}{N-4}}$. Then the functions $(u,v)$ in the Nehari manifold $N_{\omega,c}$ satisfy
$S_{\omega,c}(u,v)<S_{\omega,c}(0,Q)=S_{\omega,0}(0,Q),$
where $Q$ is the positive ground state of (1.13). In particular, $(0,Q)$ is not a minimizer of $S_{\omega,c}$ in the complex space and $(u,v)$ is a non-semitrivial solution ($u\neq 0$ and $v\neq 0$).
Proof.
It suffices to prove that the existence of $\alpha_{1},\alpha_{2}\in\mathbb{R}$ and the complex function $h\in H^{1}$ such that $(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)\in\mathcal{N}_{\omega,c}$ and $S_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)<S_{\omega,c}(0,Q)$. In fact, $(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)\in\mathcal{N}_{\omega,c}$ if and only if $N_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=0$. Since
$N_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=2Q_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)-4D(\alpha_{1}\alpha_{2}h,\alpha_{1}Q),$
by taking $\alpha_{1}\in\mathbb{R}$ satisfying
$\alpha_{1}^{2}=\frac{Q_{\omega,c}(\alpha_{2}h,Q)}{2D(\alpha_{2}h,Q)}.$ (2.11)
We see that $N_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=0$. From this point, we take $\alpha_{1}$ as in (2.11).
Now, from the identity
$Q_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=2D(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)$
and (2.11), we have
$S_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)=$ $Q_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)-D(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)$
$=$ $\frac{1}{2}Q_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)$
$=$ $\frac{\alpha_{1}^{2}}{2}Q_{\omega,c}(\alpha_{2}h,Q)$
$=$ $\frac{Q_{\omega,c}^{2}(\alpha_{2}h,Q)}{4D(\alpha_{2}h,Q)}.$
Notice that
$S_{\omega,c}(0,Q)=$ $\frac{1}{2}\int|\nabla Q|^{2}+\frac{3\gamma\omega}{2}\|Q\|_{L^{2}}^{2}+\frac{1}{2}c\cdot P(0,Q)-\frac{9}{4}\int|Q|^{4}$
$=$ $\frac{1}{2}\int|\nabla Q|^{2}+\frac{3\gamma\omega}{2}\|Q\|_{L^{2}}^{2}-\frac{9}{4}\int|Q|^{4}$
$=$ $\frac{1}{4}Q_{\omega,0}(0,Q),$
since $Q$ is a real function and $P(0,Q)=0$. Thus $S_{\omega,c}(\alpha_{1}\alpha_{2}h,\alpha_{1}Q)<S_{\omega,0}(0,Q)$ is equivalent to
$Q_{\omega,c}^{2}(\alpha_{2}h,Q)<4D(\alpha_{2}h,Q)Q_{\omega,0}(0,Q).$ (2.12)
Both sides of (2.12) are polynomials of degree $4$ in $\alpha_{2}$. The leading coefficient of the polynomial in the left-hand side is
$\left(\frac{1}{2}\int|\nabla h|^{2}+\frac{\omega}{2}|h|^{2}\right)^{2}$
whereas the leading coefficient of the polynomial in the right-hand side is
$\frac{1}{9}\int|h|^{4}\left(\frac{1}{2}\int|\nabla Q|^{2}+\frac{3\gamma\omega}{2}|Q|^{2}\right).$
Therefore, (2.12) holds for $\alpha_{2}$ sufficiently large, provided that
$Q_{\omega,c}^{2}(h,0)<\frac{1}{9}\|h\|_{L^{4}}^{4}Q_{\omega,0}(0,Q).$ (2.13)
So, it will show that (2.13) holds for some $h\in H^{1}$. For that, assume that $h(x)=Q(\lambda x)$ for some $\lambda\in\mathbb{R}$ to be determined. With this definition, (2.13) is equivalent to
$\left(\frac{\lambda^{2}}{2}\int|\nabla Q|^{2}+\frac{\omega}{2}\int|Q|^{2}\right)^{2}<\frac{\lambda^{N}}{9}Q_{\omega,0}(0,Q)\|Q\|_{L^{4}}^{4}.$ (2.14)
From the Pohozaev identities (2.4) and (2.1), we have
$\int|\nabla Q|^{2}=\frac{3\gamma\omega N}{4-N}\int|Q|^{2}=\frac{9N}{4}\int|Q|^{4}.$
Injecting this into (2.14), we obtain
$\left(\frac{\lambda^{2}}{2}\frac{3\gamma\omega N}{4-N}\int|Q|^{2}+\frac{\omega}{2}\int|Q|^{2}\right)^{2}<\frac{\lambda^{N}}{9}\left(\frac{1}{2}\frac{3\gamma\omega N}{4-N}\int|Q|^{2}+\frac{3\gamma\omega}{2}\int|Q|^{2}\right)\frac{12\gamma\omega}{9(4-N)}\int|Q|^{2}.$ (2.15)
That is
$\left(\frac{\lambda^{2}}{2}\frac{3\gamma\omega N}{4-N}+\frac{\omega}{2}\right)^{2}<\frac{\lambda^{N}}{9}\left(\frac{1}{2}\frac{3\gamma\omega N}{4-N}+\frac{3\gamma\omega}{2}\right)\frac{12\gamma\omega}{9(4-N)}.$
That is
$\left(\lambda^{2}3\gamma N+(4-N)\right)^{2}<\frac{32\gamma^{2}\lambda^{N}}{9}.$
Let
$f(\lambda)=\left(\lambda^{2}3\gamma N+(4-N)\right)-\left(\frac{32\gamma^{2}\lambda^{N}}{9}\right)^{\frac{1}{2}}.$
It is easy to see that $f$ has a global minimum at the point $\lambda_{0}=\left(\frac{9}{\sqrt{2}}\right)^{\frac{2}{N-4}}$. In addition, $f(\lambda_{0})=(4-N)(1-3\gamma\lambda_{0}^{2})$. By assumption, $f(\lambda_{0})<0$, which means that $01-3\gamma\lambda_{0}^{2}<$ 0. If we choose that $3\gamma>\left(\frac{9}{\sqrt{2}}\right)^{\frac{4}{N-4}}$, then we see that (2.15) holds for $\lambda=\lambda_{0}$ and the proof of this Lemma is complete. ∎
Proof of Theorem 1.1(ii)..
From Lemma 2.11, one can obtain the Theorem 1.1 (ii). ∎
Next, we give the proof of Theorem 1.2.
Proof of Theorem 1.2..
Assume that $(u_{\omega},v_{\omega})$ be the solution of system (1.9). Then, from the Pohozaev identities (2.4) and (2.1), we deduce that
$(4-N)\int(|\nabla u_{\omega}|^{2}+|\nabla v_{\omega}|^{2})=N\omega(\|u_{\omega}\|_{L^{2}}^{2}+3\gamma\|v_{\omega}\|_{L^{2}}^{2})+(N-2)c\cdot P(u_{\omega},v_{\omega}).$ (2.16)
Case 1: $N=2$. In this case, by the identity (2.16), we have
$\int(|\nabla u_{\omega}|^{2}+|\nabla v_{\omega}|^{2})=\omega(\|u_{\omega}\|_{L^{2}}^{2}+3\gamma\|v_{\omega}\|_{L^{2}}^{2}).$
If $\omega\leq 0$, it follows that
$\int(|\nabla u_{\omega}|^{2}+|\nabla v_{\omega}|^{2})\leq 0,$
then it must be $(u_{\omega},v_{\omega})=(0,0)$. This implies that Theorem 1.2 (i) holds.
Case 2: $N=1$ or $N=3$. Notice that in this case,
$|(N-2)c\cdot P(u_{\omega},v_{\omega})|$ $\leq|c||P(u_{\omega},v_{\omega})|$
$\leq|c|\big{(}\|u\|_{L^{2}}\|\nabla u\|_{L^{2}}+\gamma\|v\|_{L^{2}}\|\nabla v\|_{L^{2}}\big{)}$
$\leq|c|\left[\frac{1}{2}\big{(}\|u\|^{2}_{L^{2}}+3\gamma\|v\|^{2}_{L^{2}}\big{)}+\frac{1}{2}\|\nabla u\|^{2}_{L^{2}}+\frac{\gamma}{6}\|\nabla v\|^{2}_{L^{2}}\right].$
Combining with identity (2.16), then
$\left(4-N-\frac{|c|}{2}\right)\int|\nabla u_{\omega}|^{2}+\left(4-N-\frac{\gamma|c|}{6}\right)\int|\nabla v_{\omega}|^{2}\leq\left(wN+\frac{|c|}{2}\right)(\|u_{\omega}\|_{L^{2}}^{2}+3\gamma\|v_{\omega}\|_{L^{2}}^{2}).$
Thus, if $wN+\frac{|c|}{2}\leq 0$, $04-N-\frac{|c|}{2}>$ 0 and $04-N-\frac{\gamma|c|}{6}>$ 0, then it must be $(u_{\omega},v_{\omega})=(0,0)$. Thus, Theorem 1.2 (ii) holds. ∎
3 High frequency limit
In this section, we aim to study the high frequency limit of the boosted ground state of (1.1) when the frequency is $\omega\to+\infty$.
Lemma 3.1.
[18, Theorem 2.1] Let $1\leq N\leq 3$, $\omega>0$. Then there exists at least one ground state $(u_{0},v_{0})$ to system (1.7), which is radially symmetric, $v_{0}$ is positive and $u_{0}$ is either positive or identically zero.
Now we give the following scaling property.
Lemma 3.2.
Let $(\omega,c)\in\mathbb{R}\times\mathbb{R}^{N}$ satisfy $\omega>0$ and $c\neq 0$. Then
$\mu_{\omega,c}=|c|^{4-N}\mu_{\frac{\omega}{|c|^{2}},\frac{c}{|c|}}.$
Proof.
Let $\{(u_{n},v_{n})\}$ be a minimizing sequence for $\mu_{\omega,c}$, that is,
$N_{\omega,c}(u_{n},v_{n})=0,~{}~{}S_{\omega,c}\to\mu_{\omega,c},$
where $N_{\omega,c}$ and $S_{\omega,c}$ are defined in (2.2) and (1.10), respectively.
Let $(\tilde{u}_{n},\tilde{v}_{n})=\left(\frac{1}{|c|}u_{n}\left(\frac{x}{|c|}\right),\frac{1}{|c|}v_{n}\left(\frac{x}{|c|}\right)\right)$. Then
$N_{\frac{\omega}{|c|^{2}},\frac{c}{|c|}}(\tilde{u}_{n},\tilde{v}_{n})=|c|^{N-4}N_{\omega,c}(u_{n},v_{n})=0,$
$S_{\frac{\omega}{|c|^{2}},\frac{c}{|c|}}(\tilde{u}_{n},\tilde{v}_{n})=|c|^{N-4}S_{\omega,c}\to|c|^{N-4}\mu_{\omega,c}.$
Hence, we can obtain the desired result. ∎
Let
$(\tilde{u}_{\omega},\tilde{v}_{\omega})=\frac{1}{\omega^{\frac{1}{2}}}\left(u\left(\frac{x}{\omega^{\frac{1}{2}}}\right),v\left(\frac{x}{\omega^{\frac{1}{2}}}\right)\right).$ (3.1)
From Lemma 3.2, we have
$N_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=\frac{1}{\omega^{\frac{4-N}{2}}}N_{\omega,c}(u,v),~{}~{}S_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=\frac{1}{\omega^{\frac{4-N}{2}}}S_{\omega,c}(u,v).$
This means that
$\mu_{\omega,c}=\omega^{\frac{4-N}{2}}\mu_{1,\frac{c}{\sqrt{\omega}}}.$
So, $(u,v)$ is a minimizer of $\mu_{\omega,c}$ is equivalent to the fact that $(\tilde{u},\tilde{v})$ is a minimizer of $\mu_{1,\frac{c}{\sqrt{\omega}}}$.
Now we give the following uniformly bounded lemma.
Lemma 3.3.
Let $1\leq N\leq 3$. Then $(\tilde{u}_{\omega},\tilde{v}_{\omega})$ is uniformly bounded in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$ as $|c|$ is bounded and $\omega\rightarrow+\infty$.
Proof.
Let $(u_{0},v_{0})$ be a radial minimizer of $\mu_{1,0}$ given by Lemma 3.1, then
$\frac{c}{\sqrt{\omega}}\cdot P(u_{0},v_{0})=0.$
From this fact, we deduce that
$N_{1,\frac{c}{\sqrt{\omega}}}(u_{0},v_{0})=0.$
This implies that
$\mu_{1,\frac{c}{\sqrt{\omega}}}\leq\mu_{1,0}.$ (3.2)
Hence, by the definition of $\tilde{u}_{\omega}$ and $\tilde{v}_{\omega}$ (see (3.1)), we can obtain that $(\tilde{u}_{\omega},\tilde{v}_{\omega})$ is uniformly bounded in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$ as $|c|$ is bounded and $\omega\rightarrow+\infty$. ∎
In the next lemma we will give the relation between $\mu_{1,0}$ and $\mu_{1,\frac{c}{\sqrt{\omega}}}$.
Lemma 3.4.
Let $1\leq N\leq 3$. Then $0\leq\mu_{1,0}-\mu_{1,\frac{c}{\sqrt{\omega}}}\rightarrow 0$ as $|c|$ is bounded and $\omega\rightarrow+\infty$.
Proof.
From the Pohozaev identities (2.4) and (2.1), $(\tilde{u}_{\omega},\tilde{v}_{\omega})$ satisfies
$\frac{N-2}{2}\int(|\nabla\tilde{u}_{w}|^{2}+|\nabla\tilde{v}_{w}|^{2})+\frac{N}{2}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)$
$+\frac{N-1}{2}\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})=ND(\tilde{u}_{\omega},\tilde{v}_{\omega}).$ (3.3)
Since $(\tilde{u}_{\omega},\tilde{v}_{\omega})$ is the solution of (1.9), then we have
$N_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=2Q_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})-4D(\tilde{u}_{\omega},\tilde{v}_{\omega})=\omega^{\frac{N-4}{2}}N_{\omega,c}(u,v)=0.$ (3.4)
By the definition of $N_{1,\frac{c}{\sqrt{\omega}}}$ (see (2.2)), (3.4) is equivalent to
$\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})-4D(\tilde{u}_{\omega},\tilde{v}_{\omega})=0.$ (3.5)
Combining (3) and (3.5), we get
$(4-N)\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)=N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$ (3.6)
Notice that $|c|$ is bounded, by Hölder inequality, Young inequality and Lemma 3.3, we have, as $\omega\to\infty$,
$\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\to 0$ (3.7)
and
$\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\leq\frac{1}{2}\left(\int|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}+|\tilde{u}_{\omega}|^{2}+3\gamma|\tilde{v}_{\omega}|^{2}\right).$
Hence, by the Gagliardo-Nirenberg inequality (1.16), we
$\frac{1}{2}\left(\int|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}+|\tilde{u}_{\omega}|^{2}+3\gamma|\tilde{v}_{\omega}|^{2}\right)$
$<$ $2Q_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=4D(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$\leq$ $4C_{opt}^{(1)}\left(\int|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}+|\tilde{u}_{\omega}|^{2}+3\gamma|\tilde{v}_{\omega}|^{2}\right)^{2}.$
This means that
$\left(\int|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}+|\tilde{u}_{\omega}|^{2}+3\gamma|\tilde{v}_{\omega}|^{2}\right)\gtrsim 1,$
for $\omega$ is sufficiently large. Combining (3.6) and (3.7), then
$\left(\int|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)\gtrsim 1,\ \ \left(\int|\tilde{u}_{\omega}|^{2}+3\gamma|\tilde{v}_{\omega}|^{2}\right)\gtrsim 1.$ (3.8)
Let
$\left(\tilde{u}^{\lambda}_{\omega},\tilde{v}^{\lambda}_{\omega}\right)=\left(\tilde{u}_{\omega}(x/\lambda),\tilde{v}_{\omega}(x/\lambda)\right),~{}~{}\lambda>0.$ (3.9)
Then
$N_{1,0}\left(\tilde{u}^{\lambda}_{\omega},\tilde{v}^{\lambda}_{\omega}\right)=\lambda^{N-2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\lambda^{N}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-4\lambda^{N}D(\tilde{u}_{\omega},\tilde{v}_{\omega}),$
where $N_{1,0}$ is defined in (2.2) with $\omega=1$ and $c=0$.
From (3) and (3.6), we have
$N_{1,0}\left(\tilde{u}^{\lambda}_{\omega},\tilde{v}^{\lambda}_{\omega}\right)$
$=$ $\lambda^{N-2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)-\frac{\lambda^{N}}{4-N}\left(N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\right)$
$=$ $\frac{\lambda^{N-2}}{4-N}\left(N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\right)$
$-\frac{\lambda^{N}}{4-N}\left(N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\right).$ (3.10)
Notice that
$N_{1,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})=N_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})-\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})=-\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$
Case 1. If $\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0$, then $N_{1,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})<0$. Obviously, using (3.8), if $|c|$ is bounded and $\omega\to+\infty$,
$N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0,$
$N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0,$
and
$(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})<\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$
Hence, from (3), there exists $\lambda_{\omega}\in(0,1)$ such that
$N_{1,0}\left(\tilde{u}^{\lambda_{\omega}}_{\omega},\tilde{v}^{\lambda_{\omega}}_{\omega}\right)=0.$
Case 2. If $\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})<0$, then $N_{1,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})>0$. Similarly,
$N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0,$
$N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0,$
Since $1\leq N\leq 3$, then
$(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$
Hence, from (3) there exists $\lambda_{\omega}\in(1,+\infty)$ such that
$N_{1,0}\left(\tilde{u}^{\lambda_{\omega}}_{\omega},\tilde{v}^{\lambda_{\omega}}_{\omega}\right)=0.$
We claim that
$\lambda_{\omega}\in(1,+\infty)~{}~{}\text{uniformly bounded}.$
In fact, by assumption, $N\leq 3$, $|c|$ is bounded and $\omega\to+\infty$, and the estimates (3.7) and (3.8), we deduce that
$\frac{1}{2}\left(N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\right)$
$<$ $N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$
Then, from (3), we can obtain that $\lambda^{2}_{\omega}<2$. Hence, the claim is true.
Furthermore, we claim that
$\lambda_{\omega}\to 1~{}~{}\text{as}~{}~{}\omega\to+\infty.$ (3.11)
Indeed, from (3), we have
$0=$ $N_{1,0}\left(\tilde{u}^{\lambda_{\omega}}_{\omega},\tilde{v}^{\lambda_{\omega}}_{\omega}\right)$
$=$ $\frac{\lambda_{\omega}^{N-2}}{4-N}\left(N(1-\lambda_{\omega}^{2})\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+[N-2-2\lambda_{\omega}^{2}]\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\right).$
By (3.7) and (3.8), we can obtain the claim (3.11).
Now, from the identity (3.4), we deduce that
$\lambda_{\omega}^{N-2}S_{1,0}\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$
$=$ $\frac{1}{2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\frac{(\lambda_{\omega})^{2}}{2}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-(\lambda_{\omega})^{2}D(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$=$ $\frac{1}{2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\lambda_{\omega}^{2}\frac{1}{4}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)$
$-\frac{\lambda_{\omega}^{2}}{4}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)-\lambda_{\omega}^{2}\frac{1}{4}\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$=$ $\left(\frac{1}{2}-\frac{\lambda_{\omega}^{2}}{4}\right)\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\lambda_{\omega}^{2}\frac{1}{4}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-\lambda_{\omega}^{2}\frac{1}{4}\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$=$ $\frac{1}{2}Q_{1,\frac{c}{\sqrt{w}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})+\frac{1}{4}(1-\lambda_{\omega}^{2})\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)$
$+\frac{1}{4}(1-\lambda_{\omega}^{2})\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-\frac{1}{4}(1+\lambda_{\omega}^{2})\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})$
[…]
where $Q_{1,\frac{c}{\sqrt{\omega}}}$ is defined by (2.1) and in the last step, we used the Lemma 3.3 and (3.11). This means that
$\mu_{1,0}\leq\mu_{1,\frac{c}{\sqrt{\omega}}}+o(1),~{}~{}\text{as}~{}~{}\omega\to+\infty.$ (3.12)
Combining (3.2) and (3.12), we can obtain the desired result. Now we have completed the proof of this Lemma. ∎
Lemma 3.5.
Up to a subsequence, $(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})\to(\tilde{u}_{\infty},\tilde{v}_{\infty})$ strongly in $H^{1}$ as $\omega\rightarrow+\infty$. Moreover, $(\tilde{u}_{\infty},\tilde{v}_{\infty})$ is a minimizer sequence of $\mu_{1,0}$.
Proof.
From Lemma 3.4, it is known that
$N_{1,0}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})=0$
and
$S_{1,0}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})\to\mu_{1,0}~{}~{}\text{as}~{}~{}\omega\to\infty.$
This means that $\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$ is the minimizing sequence of (2.3). By the definition of $(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})$ (see (3.9)) and Lemma 3.3, we get $(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{u}_{\omega}^{\lambda_{\omega}})$ is also uniformly bounded in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$.
By similar argument as Lemma 2.6 and 2.9, there exist $\{y_{\omega}\}$, $(\tilde{u}_{\infty},\tilde{v}_{\infty})\in H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})\backslash\{(0,0)\}$, and a subsequence $\tau_{y_{\omega}}\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$ (still denoted by $\tau_{y_{\omega}}\left(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}\right)$ such that
$\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})\rightharpoonup(u_{\infty},v_{\infty})~{}~{}\text{weakly in}~{}~{}H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N}).$
By weakly convergence of $\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})$ and Lemma 2.8, we have
$Q_{1,0}\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-Q_{1,0}\left(\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-(\tilde{u}_{\infty},\tilde{v}_{\infty})\right)\to Q_{1,0}(\tilde{u}_{\infty},\tilde{v}_{\infty}).$
Since $Q_{1,0}(\tilde{u}_{\infty},\tilde{v}_{\infty})>0$, up to a subsequence, we have
$\frac{1}{2}\lim_{\omega\to\infty}Q_{1,0}\left(\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-(\tilde{u}_{\infty},\tilde{v}_{\infty})\right)<\frac{1}{2}\lim_{\omega\to\infty}Q_{1,0}(\tau_{\omega}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}))=\mu_{1,0}.$
By a similar argument as Lemma 2.7, we can obtain that
$\mu_{1,0}\leq\frac{1}{2}Q_{1,0}.$
Then, using the above and the weakly lower semi-continuity of norms, we obtain
$\mu_{1,0}\leq\frac{1}{2}Q_{1,0}(\tilde{u}_{\infty},\tilde{v}_{\infty})\leq\frac{1}{2}\lim_{\omega\to\infty}Q_{1,0}(\tau_{\omega}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}}))=\mu_{1,0}.$
Hence, we have
$Q_{1,0}\left(\tau_{y_{\omega}}(\tilde{u}_{\omega}^{\lambda_{\omega}},\tilde{v}_{\omega}^{\lambda_{\omega}})-(\tilde{u}_{\infty},\tilde{v}_{\infty})\right)\to 0,$
which implies that
$(\tilde{u}_{\omega}^{t_{\omega}},\tilde{u}_{\omega}^{t_{\omega}})\to(\tilde{u}_{\infty},\tilde{v}_{\infty})~{}~{}\text{strongly in}~{}~{}H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N}).$
Now we complete the proof of Lemma 3.5. ∎
Proof of Theorem 1.4..
By the definitions of (3.1) and (3.9) and Lemma 3.5, one can easily obtain
$\frac{1}{\omega^{\frac{1}{2}}}\left(u_{\omega}\left(\frac{x}{\lambda_{\omega}\omega^{\frac{1}{2}}}\right),v_{\omega}\left(\frac{x}{\lambda_{\omega}\omega^{\frac{1}{2}}}\right)\right)\to(u_{\infty},v_{\infty})~{}~{}\text{in}~{}~{}H^{1}(\mathbb{R}^{N}),$
where $\lambda_{\omega}\to 1$ (see (3.11)).
On the other hand, by the weak convergence and the norm convergence, we have
$\frac{1}{\omega^{\frac{1}{2}}}\left(u_{\omega}\left(\frac{x}{\lambda_{\omega}\omega^{\frac{1}{2}}}\right),v_{\omega}\left(\frac{x}{\lambda_{\omega}\omega^{\frac{1}{2}}}\right)\right)\to\frac{1}{\omega^{\frac{1}{2}}}\left(u_{\omega}\left(\frac{x}{\omega^{\frac{1}{2}}}\right),v_{\omega}\left(\frac{x}{\omega^{\frac{1}{2}}}\right)\right)~{}~{}\text{in}~{}~{}H^{1}(\mathbb{R}^{N}),$
as $\lambda_{\omega}\to 1$. Hence, combining the above two relations, we conclude the proof of Theorem 1.4. ∎
4 Global existence
In this section, we aim to obtain the global existence. First, we introduce a subset of the energy space
$A^{+}_{\omega,c}=\{(u,v)\in H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})~{}:~{}S_{\omega,c}(u,v)\leq\mu_{\omega,c},~{}N_{\omega,c}(u,v)\geq 0\}.$ (4.1)
Now we show that $A^{+}_{\omega,c}$ is an invariant set under the flow.
Lemma 4.1.
Assume (1.11). Then the set $A^{+}_{\omega,c}$ is invariant under the flow of (1.1).
Proof.
By the similar argument as [9, 12], we can easily obtain this result. Here we omit it. ∎
Lemma 4.2.
Assume the condition (1.11) holds. If the initial data $(u_{0},v_{0})\in A^{+}_{\omega,c}$, then the $H^{1}$ solution $(u(t),v(t))$ of (1.1) exists globally in time and
$\sup_{t\in\mathbb{R}}\|(u(t),v(t))\|_{H^{1}\times H^{1}}\leq C(\|(u_{0},v_{0})\|_{H^{1}\times H^{1}})<\infty.$
Proof.
Combining Lemma 4.1 and the definitions of $S_{\omega,c}$, $Q_{\omega,c}$ and $N_{\omega,c}$, we have
$\mu_{\omega,c}\geq$ $S_{\omega,c}(u(t),v(t))$
$=$ $\frac{1}{2}Q_{\omega,c}(u(t),v(t))+\frac{1}{4}N_{\omega,c}(u(t),v(t))$
$\geq$ $\frac{1}{2}Q_{\omega,c}(u(t),v(t))$
$\geq$ $\frac{1}{4}\left\|\nabla u(t)-\frac{i}{2}cu(t)\right\|_{L^{2}}^{2}+\frac{1}{4}\left\|\nabla v(t)-\frac{i\gamma}{2}cv(t)\right\|_{L^{2}}^{2}.$
By the conservation of mass, we deduce
$\|\nabla u(t)\|_{L^{2}}^{2}+\|\nabla v(t)\|_{L^{2}}^{2}$
$\leq$ $C\left(\left\|\nabla u(t)-\frac{i}{2}cu(t)\right\|_{L^{2}}^{2}+\left\|\nabla v(t)-\frac{i\gamma}{2}cv(t)\right\|_{L^{2}}^{2}\right)+C\left(\|u(t)\|_{L^{2}}^{2}+\|v(t)\|_{L^{2}}^{2}\right)$
$\leq$ $C\mu_{\omega,c}+CM_{3\gamma}(u_{0},v_{0}).$
This means that the $H^{1}$ solution is globally in time. Now we complete the proof of this lemma. ∎
Proof of Theorem 1.5.
To prove (i). Let $A_{0}$ be chosen later. We claim that if $(u_{0},v_{0})\in H^{1}(\mathbb{R}^{2})$ satisfies $\|v_{0}\|_{L^{2}}^{2}<A_{0}$ and $0<\epsilon\ll\frac{1}{|c|^{2}}$ , then we claim that
$S_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)<\mu_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c},$ (4.2)
$N_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\geq 0.$ (4.3)
for large $|c|$. If this claim is true, then we can obtain
$\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\in A^{+}_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}$
for $|c|$ large enough. Hence, by Lemma 4.2, we can obtain the desired result.
Now we prove the claim. Indeed, by the definition of $Q_{\omega,c}$, we have
$Q_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}(u,v)$
$=$ $\frac{1}{2}\int|\nabla u|^{2}+\frac{1}{2}\int|\nabla v|^{2}+\frac{(1+\epsilon)|c|^{2}}{8}\|u\|_{L^{2}}^{2}+\frac{3\gamma(1+\epsilon)|c|^{2}}{8}\|v\|_{L^{2}}^{2}+\frac{1}{2}c\cdot P(u,c)$
$=$ $\frac{1}{2}\left\|\nabla\left(e^{-\frac{i}{2}c\cdot x}u\right)\right\|_{L^{2}}^{2}+\frac{1}{2}\left\|\nabla\left(e^{-\frac{i\gamma}{2}c\cdot x}v\right)\right\|_{L^{2}}^{2}$
$+\frac{\epsilon|c|^{2}}{8}\|u\|_{L^{2}}^{2}+\frac{3(1+\epsilon)\gamma|c|^{2}-\gamma^{2}|c|^{2}}{8}\|v\|_{L^{2}}^{2}.$
From Lemma 3.2 and above, (4.2) is equivalent to
$S_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
$=$ $Q_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)-D\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
$=$ $\frac{1}{2}\left\|\nabla u_{0}\right\|_{L^{2}}^{2}+\frac{1}{2}\left\|\nabla v_{0}\right\|_{L^{2}}^{2}$
$+\frac{\epsilon|c|^{2}}{8}\|u_{0}\|_{L^{2}}^{2}+\frac{3(1+\epsilon)\gamma|c|^{2}-\gamma^{2}|c|^{2}}{8}\|v_{0}\|_{L^{2}}^{2}$
$-\int\frac{1}{36}|u_{0}|^{4}+\frac{9}{4}|v_{0}|^{4}+|u_{0}|^{2}|v_{0}|^{2}+\frac{1}{9}\Re\left(e^{\frac{i(\gamma-3)}{2}c\cdot x}\bar{u}_{0}^{3}(x)v_{0}(x)\right)$
$\leq$ $|c|^{2}\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|}}.$
That is,
$\frac{1}{2}\left\|\nabla u_{0}\right\|_{L^{2}}^{2}+\frac{1}{2}\left\|\nabla v_{0}\right\|_{L^{2}}^{2}+\frac{\epsilon|c|^{2}}{8}\|u_{0}\|_{L^{2}}^{2}$
$-\int\frac{1}{36}|u_{0}|^{4}+\frac{9}{4}|v_{0}|^{4}+|u_{0}|^{2}|v_{0}|^{2}+\frac{1}{9}\Re\left(e^{\frac{i(\gamma-3)}{2}c\cdot x}\bar{u}_{0}^{3}(x)v_{0}(x)\right)$
$\leq$ $|c|^{2}\left(\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|}}-\frac{\gamma(3(1+\epsilon)-\gamma)}{8}\|v_{0}\|_{L^{2}}^{2}\right).$ (4.4)
By the Gagliardo-Nirenberg inequality, we have
$\frac{1}{2}\left\|\nabla u_{0}\right\|_{L^{2}}^{2}+\frac{1}{2}\left\|\nabla v_{0}\right\|_{L^{2}}^{2}-\int\frac{1}{36}|u_{0}|^{4}+\frac{9}{4}|v_{0}|^{4}+|u_{0}|^{2}|v_{0}|^{2}$
$\geq$ $\frac{1}{2}\left\|\nabla u_{0}\right\|_{L^{2}}^{2}+\frac{1}{2}\left\|\nabla v_{0}\right\|_{L^{2}}^{2}-\frac{1}{2}\frac{M_{3\gamma}(u_{0},v_{0})}{M_{3\gamma}(Q^{*},P^{*})}(\left\|\nabla u_{0}\right\|_{L^{2}}^{2}+\left\|\nabla v_{0}\right\|_{L^{2}}^{2})$
$=$ $\frac{1}{2}\left(1-\frac{M_{3\gamma}(u_{0},v_{0})}{M_{3\gamma}(Q^{*},P^{*})}\right)(\left\|\nabla u_{0}\right\|_{L^{2}}^{2}+\left\|\nabla v_{0}\right\|_{L^{2}}^{2})>0,$
where in the last step we used the assumption $M_{3\gamma}(u_{0},v_{0})<M_{3\gamma}(Q^{*},P^{*})$, (see (1.19)).
Let
$A_{0}=\frac{8}{\gamma(3(1+\epsilon)-\gamma)}\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|}}.$
Then, by $\gamma<3$ and Lemma 2.5, we get $A_{0}>0$. Let $\|v_{0}\|_{L^{2}}^{2}<A_{0}$, then
$|c|^{2}\left(\mu_{\frac{1+\epsilon}{4},\frac{c}{|c|}}-\frac{\gamma(3(1+\epsilon)-\gamma)}{8}\|v_{0}\|_{L^{2}}^{2}\right)\to\infty~{}~{}\text{as}~{}~{}|c|\to\infty.$
On the other hand, the Riemann-Lebesgue theorem implies that
$\frac{1}{9}\int\Re\left(e^{\frac{i(\gamma-3)}{2}c\cdot x}\bar{u}_{0}^{3}(x)v_{0}(x)\right)\to 0~{}~{}\text{as}~{}~{}|c|\to\infty.$ (4.5)
From (4.5), we deduce that the left-hand side of (4) is bounded as $|c|$ is sufficiently large, since $0<\epsilon\ll\frac{1}{|c|^{2}}$. Therefore, (4) holds for $|c|$ large enough, and so does (4.2). In addition, by (4.5), we obtain
$N_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
$=$ $2Q_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)-4D\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)$
$=$ $\left\|\nabla u_{0}\right\|_{L^{2}}^{2}+\left\|\nabla v_{0}\right\|_{L^{2}}^{2}+\frac{\epsilon|c|^{2}}{4}\|u_{0}\|_{L^{2}}^{2}+\frac{3(1+\epsilon)\gamma|c|^{2}-\gamma^{2}|c|^{2}}{4}\|v_{0}\|_{L^{2}}^{2}$
$-4\left(\int\frac{1}{36}|u_{0}|^{4}+\frac{9}{4}|v_{0}|^{4}+|u_{0}|^{2}|v_{0}|^{2}+\frac{1}{9}\Re\left(e^{\frac{i(\gamma-3)}{2}c\cdot x}\bar{u}_{0}^{3}(x)v_{0}(x)\right)\right).$
Since $u_{0},v_{0}\in H^{1}$, then there exists a constant $C_{0}>0$ such that $\|u_{0}\|_{L^{4}},\|v_{0}\|_{L^{4}}\leq C_{0}$. Hence, for large $|c|$ and $30<\gamma<$ 3,
$\frac{\gamma|c|^{2}(3(1+\epsilon)-\gamma)}{4}\|v_{0}\|_{L^{2}}^{2}-4\int\frac{1}{36}|u_{0}|^{4}+\frac{9}{4}|v_{0}|^{4}+|u_{0}|^{2}|v_{0}|^{2}\geq 0$
holds. Combining the above estimates, we can obtain that
$N_{\frac{|\sqrt{1+\epsilon}c|^{2}}{4},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\geq 0$
for $|c|$ large enough. Then, (4.3) holds. Hence, (i) holds.
To prove (ii). Let $0<\epsilon\ll\frac{1}{|c|^{2}}$ and
$B_{0}=\frac{24}{(1+\epsilon)\gamma-3}\mu_{\frac{(1+\epsilon)\gamma}{12},\frac{c}{|c|}}.$
If the initial data $(u_{0},v_{0})\in H^{1}(\mathbb{R}^{2})\times H^{1}(\mathbb{R}^{2})\backslash\{(0,0)\}$ satisfies $\|u_{0}\|_{L^{2}}^{2}<B_{0}$, then we can obtain, for $|c|$ large enough,
$S_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)<\mu_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c},$
$N_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c}\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\geq 0.$
This means that, for $|c|$ large enough,
$\left(e^{\frac{i}{2}c\cdot x}u_{0},e^{\frac{i\gamma}{2}c\cdot x}v_{0}\right)\in A^{+}_{\frac{(1+\epsilon)\gamma|c|^{2}}{12},c}.$
Hence, (ii) holds.
Now we complete the proof of Theorem 1.5. ∎
Appendix A Appendix
Proof of Lemma 2.1..
Multiplying the first equation in (1.9) by $\bar{\phi}$, the second one by $\bar{\psi}$, integrating over $\mathbb{R}^{N}$ and using integration by parts, we have
$\int|\nabla\phi|^{2}+\omega\int|\phi|^{2}+(ic\cdot\nabla\phi,\phi)-\frac{1}{9}\int|\phi|^{2}-2\int|\psi|^{2}|\phi|^{2}-\frac{1}{3}\int\bar{\phi}^{3}\psi=0,$
$\int|\nabla\psi|^{2}+3\gamma\omega\int|\psi|^{2}+(i\gamma c\cdot\nabla\psi,\psi)-9\int|\psi|^{2}-2\int|\psi|^{2}|\phi|^{2}-\frac{1}{9}\int\phi^{3}\bar{\psi}=0.$
Summing the above identities, we can obtain
$\int|\nabla\phi|^{2}+\omega\int|\phi|^{2}+(ic\cdot\nabla\phi,\phi)+\int|\nabla\psi|^{2}+3\gamma\omega\int|\psi|^{2}+(i\gamma c\cdot\nabla\psi,\psi)$
$=$ $\int\left(\frac{1}{9}|\phi|^{4}+9|\psi|^{2}+4\int|\psi|^{2}|\phi|^{2}+\frac{4}{9}\int\phi^{3}\bar{\psi}\right).$
Since
$(x\cdot\nabla u,ic\cdot\nabla u)=-\frac{N-1}{2}(u,ic\cdot\nabla u),~{}~{}(x\cdot\nabla u,-\Delta u)=-\frac{N-2}{2}\int|\nabla u|^{2},$
$(x\cdot\nabla u,u)=-\frac{N}{2}\int|u|^{2},~{}~{}(x\cdot\nabla u,|u|^{p-1}u)=-\frac{N}{p+1}\int|u|^{p+1}.$
Multiplying the first equation in (1.9) by $x\cdot\nabla\bar{\phi}$, the second one by $x\cdot\nabla\bar{\psi}$, integrating over $\mathbb{R}^{N}$, using integration by parts and taking the real part, we deduce that
$-\frac{N-2}{2}\int|\nabla\phi|^{2}-\omega\frac{N}{2}\int|\phi|^{2}-\frac{N-1}{2}(\phi,ic\cdot\nabla\phi)$
$=-\frac{N}{36}\int|\phi|^{4}+2\Re\int|\psi|^{2}\phi x\cdot\nabla\bar{\phi}+\frac{1}{3}\Re\int\bar{\phi}^{2}\psi x\cdot\nabla\bar{\phi},$
and
$-\frac{N-2}{2}\int|\nabla\psi|^{2}-3\gamma\omega\frac{N}{2}\int|\psi|^{2}-\gamma\frac{N-1}{2}(\psi,ic\cdot\nabla\psi)$
$=$ $-\frac{9N}{4}\int|\psi|^{4}+2\Re\int|\phi|^{2}\psi x\cdot\nabla\bar{\psi}+\frac{1}{9}\Re\int{\phi}^{3}x\cdot\nabla\bar{\psi}$
$=$ $-\frac{9N}{4}\int|\psi|^{4}-2\Re\int|\psi|^{2}\phi x\cdot\nabla\bar{\phi}-N\int|\psi|^{2}|\phi|^{2}-\frac{1}{3}\Re\int\phi^{2}\bar{\psi}x\cdot\nabla\phi-\frac{N}{9}\Re\int\phi^{3}\bar{\psi}.$
From above two identities, we obtain
$-\frac{N-2}{2}\int\left(|\nabla\phi|^{2}+|\nabla\psi|^{2}\right)-\frac{N}{2}\int\left(\omega|\phi|^{2}+3\gamma\omega|\psi|^{2}\right)$
$-\frac{N-1}{2}\left[(\phi,ic\cdot\nabla\phi)+\gamma(\psi,ic\cdot\nabla\psi)\right]$
$=$ $-\frac{N}{36}\int|\phi|^{4}-\frac{9N}{4}\int|\psi|^{4}-N\int|\psi|^{2}|\phi|^{2}-\frac{N}{9}\Re\int\phi^{3}\bar{\psi}.$
This completes the proof of Lemma 2.1. ∎
Appendix B Appendix
In this section, we aim to prove the sharp Gagliardo-Nirenberg-type inequality (1.17). Now, we give the following lemma.
Lemma B.1.
Let $1\leq N\leq 3$, system (1.18) has a ground state solution $(Q^{*},P^{*})$. Moreover, the following identities hold.
$\int|\nabla Q^{*}|^{2}+\int|Q^{*}|^{2}=\frac{1}{9}\int|Q^{*}|^{4}+2\int|Q^{*}|^{2}|P^{*}|^{2},$
$\int|\nabla P^{*}|^{2}+3\gamma\int|P^{*}|^{2}=9\int|P^{*}|^{4}+2\int|Q^{*}|^{2}|P^{*}|^{2},$
$\frac{N-2}{2}\int(|\nabla Q^{*}|^{2}+|\nabla P^{*}|^{2})+\frac{N}{2}\int|Q^{*}|^{2}+\frac{3\gamma N}{2}\int|P^{*}|^{2}$
$-\frac{N}{4}\int\left(\frac{1}{9}|Q^{*}|^{4}+9|P^{*}|^{4}+4|Q^{*}|^{2}|P^{*}|^{2}\right)=0.$
Proof.
By the similar argument as [18], we can obtain this result. Here we omit it. ∎
Proof of inequality (1.17)..
By the Gagliardo-Nirenberg inequality
$\|f\|_{L^{4}}^{4}\leq C\|\nabla f\|_{L^{2}}^{2}\|f\|_{L^{2}}^{2},$
we have
$\int\left(\frac{1}{36}|u|^{4}+\frac{9}{4}|v|^{4}+u^{2}v^{2}\right)\leq CK(u,v)M_{3\gamma}(u,v).$
Now we define
$J(u,v)=\frac{K(u,v)M_{3\gamma}(u,v)}{\int\left(\frac{1}{36}|u|^{4}+\frac{9}{4}|v|^{4}+u^{2}v^{2}\right)}.$
The infimum of $J(u,v)$ is attained at a pair of real function $(Q^{*},P^{*})$, that is
$\inf J(u,v)=J(Q^{*},P^{*}),$
if and only if, up to scaling, $(Q^{*},P^{*})$ is a positive ground state solution of (1.18). In fact, this is true. A similar argument can be found in [18].
Hence, we can obtain that
$\frac{1}{C_{opt}^{(2)}}=J(Q^{*},P^{*})=2M_{3\gamma}(Q^{*},P^{*}).$
Acknowledgments.
This work was supported by the National Natural Science Foundation of China (12301090 and 11801519) and Fundamental Research Program of Shanxi Province NO.202203021222126. Part of this work was done by Y.L. during his postdoctoral studies at Central China Normal University, and he would like to thank the professor Shuangjie Peng for fruitful discussions and constant encouragement.
Data Availability
We do not analyse or generate any datasets, because our work proceeds within a theoretical and mathematical approach.
Conflict of interest
The authors declare no potential conflict of interest.